OpenStudy (anonymous):

How do you show that a set of vectors from the basis of R^3?

OpenStudy (anonymous):

The questions specifically asks about v1=[-1, 0, 3] v2 = [-1, 1, 0] and v3= [0, 1, 2]

OpenStudy (anonymous):

I know that it must be a spanning set and linearly independent and I know how to show it is linearly independent, I just don't know how to show it is a spanning set.

OpenStudy (anonymous):

aahh

OpenStudy (anonymous):

Sorry, that is above my head, we haven't done determinants

OpenStudy (anonymous):

Take any point in R3 say (15, 14,-3) k1v1+k2v2+k3v3 = (15,14,-3) k1[-1, 0, 3]+ k2[-1, 1, 0] + k3[0, 1, 2] (-k1-k2,k2+k3,3k1+ 2k3) = ( 15,14,-3) solve for k1 k2 k3 you must get them

OpenStudy (anonymous):

OK I think I get that

OpenStudy (anonymous):

k1 k2 k3 are scalars ?